### v5

# Distribution System Reconfiguration to Mitigate Load Altering Attacks via Stackelberg Games

Abstract
The integration of IoT-controllable devices in power systems (such as smart electric vehicle charging stations, heat pumps, etc.), despite their benefits, raises novel cybersecurity concerns. Vulnerabilities in these devices can be leveraged to launch load-altering attacks (LAAs) that can potentially compromise the safety of power systems. In this paper, we analyze the impact of LAAs on the voltage profile of distribution networks (DNs). We first derive closed-form expressions to quantify the attacks’ impact. Using the insights derived from this analysis, we then propose a reactive defense method to mitigate LAAs based on reconfiguring the DNs. We also study optimal defense strategies that are robust to LAAs by exploiting non-cooperative sequential game theory. The proposed solution takes into account the potential uncertainties in the attack localization. Furthermore, we propose a Bayesian optimization (BO) approach to compute the Stackelberg equilibrium of the game, which reduces the computational burden. Our results show that attacks launched on the deepest nodes in the DN have the most detrimental effect on the grid voltage profile. Furthermore, the proposed game-theoretic strategy successfully mitigates the effect of the attack while ensuring minimum system reconfiguration.
Index Terms:
Distribution network, Cybersecurity, Load-altering attack (LAA), Reconfiguration, Stackelberg game, Bayesian optimization.
I Introduction
Internet-of-Things (IoT) enabled devices offer enhanced end-user convenience, improved efficiency and flexibility to power systems for load peak management, which has driven a notable surge in their adoption. However, beyond their evident benefits, these devices also present potential vulnerabilities, serving as entry points for cyber attackers to exploit and compromise the security of power systems. Specifically, load-altering attacks (LAA) in power networks with high IoT-enabled device penetration pose a significant cybersecurity threat [2, 3, 4, 5].
The concept of LAAs was first introduced in [3] in which adversaries turn a group of IoT-controllable electrical loads into bots and turn them on/off simultaneously to harm the stability of the system. The manipulation of loads disrupts the balance between power generation and demand, leading to frequency instabilities in transmission networks [6, 4]. In distribution networks (DNs), LAA can result in elevated line flows, causing higher voltage drops leading to voltage constraint violation [7]. Furthermore, the data required to launch successful LAAs can be obtained from publicly available information [8] or circumvented [9].
I-A Literature Survey
LAAs have gained significant interest over the last few years. We divide the existing works into two groups: 1) attack impact analysis and viability; and 2) attack mitigation.
Attack Impact Analysis and Viability: Researchers in [4] investigated the impact of LAAs on transmission systems and identified several effects, including line failures, unsafe frequency deviations, disruption in grid restarting, and increased operational costs. The research in [10] examined the effects of LAA under a more realistic setting consisting of protection (such as $N-1$ security) and load-shedding schemes and showed that LAA can still cause outages and islands. [11] examined the feasibility and impact of LAAs under low demand and low inertia conditions experienced during the COVID-19 pandemic, and showed that under these conditions, it can be easier to launch LAAs that lead to unsafe grid conditions. A study on the impact of attacks on IoT-enabled devices on interconnected transmission and distribution systems was carried out in [12].
Reference [2] proposed the so-called dynamic LAA (DLAA), in which the adversary continuously toggles the compromised load devices on and off, guided by a feedback control loop in response to the system’s frequency fluctuations. In [5], an analytical framework was introduced to analyze the impact of LAA on transmission systems and identify the nodes from which an attacker can launch the most effective attacks using the theory of second-order dynamical systems. In [13], a rare-event sampling algorithm was proposed that uncovers the spatial and temporal distribution of impactful DLAAs while considering the security constraint of $N-1$.
The growing popularity of electric vehicles (EVs) and the spread of EV charging stations (EVCS) make them a potential target for the LAA. Reference [14] explores the weak points of EVCS, such as firmware flaws, vulnerabilities in management systems, and the security scarcity of mobile apps that allow adversaries to launch LAAs. Subsequently, to detect such attacks, they propose model-based and data-driven approaches. Furthermore, in [15] authors study the problem of launching DLAA on coordinated distributed EVCS, while considering potential communication delays when sending attack commands. Furthermore, their attack circumvents existing defensive solutions, such as traveling wave and transient energy-based methods. They employ a robust optimization method to launch a coordinated attack from multiple locations to maximize the impact of the attack.
Attack Mitigation: Another stream of research investigates the mitigation of LAAs. The existing mitigation methods can be categorized into: i) offline; or ii) online methods.
TABLE I: State of the art on LAA mitigation methods and position of our work.
Ref Impact analyses Strategic attacker Defense approach Defense type
Transmission [2] Securing loads Preventive
[16] Robust operating points Preventive
[17] Frequency droop control Reactive
[18] EV charge/discharge Reactive
[19] Load shedding Reactive
[20] Securing devices & Optimal load management Hybrid
[21] Reactive power compensation Preventive
Dist. [7] SOPs Preventive
Our work Reconfiguration Reactive
Offline methods: Offline defenses try to install preventive measures to stop the impact of LAAs. For instance, [16] proposed algorithms to determine the operating points for generators in a way to prevent line overloads caused by potential botnet-type attacks against IoT devices. Reference [2] presented a mitigation framework based on securing the most critical loads. This method found the minimum magnitude of loads needed to be protected in order to guarantee frequency stability in the event of DLAAs. [21] proposed a zero-sum Stackelberg game formulation to install reactive power compensation (RPC) to reduce the impact of the attack. While the works above focus on transmission systems, [7] introduced a mitigation approach tailored specifically for DNs. Their research focuses on identifying optimal locations for deploying soft open points (SOPs) and refining their operation to mitigate the effects of attacks on voltage deviations.
Online methods: Despite the effectiveness of the offline methods, these measures may be too costly as the preventive features must be enabled irrespective of whether an attack occurs or not (e.g., uneconomic generator operating points to cover for LAAs). Online methods, on the other hand, involve determining defensive actions to counter the effects of LAAs once the attack is launched, via reactive measures. In [17], a cyber-resilient economic dispatch method is introduced to mitigate LAAs based on altering the frequency droop control parameters of inverter-based resources to counter the destabilizing effects of LAAs. [18] proposed a framework in which electric vehicles are designed as feedback controllers that can mitigate the impact of LAA based on $H-2$ and $H-\infty$ norms. To analyze the manoeuvres of a strategic attacker initiating DLAA, [19] introduced a multi-stage game approach. In this game, the defensive actions involved load shedding, and the ultimate objective is to achieve a strategic balance between DLAA and the necessary amount of load shedding, reaching a Nash equilibrium (NE). [20] formulated the cybersecurity of the power system in both cyber and physical layers as a game-theoretic formulation.
Reconfiguration of DNs: Reconfiguration of DNs is a well-studied topic in which the topology of the networks is changed by turning existing line switches on/off [22]. The existence of switches in lines which also can be used to isolate potential faults in lines provides a paramount amount of possible configurations for the grid. This usually takes place for reducing power loss, network balancing, voltage profile rectification, enhancing network restoration, and network reliability [23, 24, 25]. The fact that the operator can perform the changes in topology swiftly, only by sending on/off commands to switches makes this capability a great fit for responding to a potential threat.
I-B Research Gap and Contributions
Based on the discussions of Section I-A and Table I it can be inferred that there is a dearth of studies for analyzing the impact of LAA on DNs. Similarly, the majority of mitigations are designed to compensate for attacks in transmission grids. This paper addresses this gap in LAA studies in DNs. Furthermore, this paper is the first to consider mitigation against strategic attackers in the context of LAAs on DNs. Drawing from the insights presented in [26, 27], we also considered noisy system measurements and uncertainties associated with attack detection/localization in our mitigation scheme.
We start with a rigorous investigation of the impact of LAAs on DNs. We integrate the ZIP load model [28] into our formulations to obtain a realistic attack impact of DNs. Our work provides analytical insights into the impact of LAAs and identifies the most vulnerable nodes of the system. The results are then utilized to guide the design of the mitigation strategy. To take a strategic attacker into account and to defend against the worst-case attacks, we model the attacker-defender interaction using a Stackelberg game framework with the attacker as the leader and the defender as the follower. This sequential game is relevant for reactive defenses, in which mitigation is triggered only when an attack is detected. To obtain the players’ payoffs, the reconfiguration problem should be solved repeatedly. We reformulate this problem as a mixed integer linear programming (MILP) optimization by using the linearized distribution flow (LinDistFlow) and the ZP approximations. The resulting formulation provides a significant computational speed-up in calculating the game’s payoffs.
Despite simplifying the optimization, obtaining the Stackelberg equilibrium remains computationally complex in cases involving large sets of player actions since it requires repeatedly solving optimization problems to compute the payoff functions (e.g., if an attacker aims to target $n_{a}$ buses from a pool of $N$ candidate victim buses, the total number of possible attacks is given by $\binom{N}{n_{a}}$, which can be very large.). To mitigate this burden, we propose using Bayesian optimization (BO). Rather than exhaustively evaluating all possible actions, this approach constructs a probabilistic model and iteratively refines the model, focusing on the most promising actions to identify optimal strategies swiftly.
This paper significantly extends our preliminary work [1] in terms of using the insights derived from our attack impact analysis to design a mitigation strategy to counter LAAs. The key contributions are summarized as follows:
$\bullet$ Deriving closed-form expressions of the bus voltages of the DN in the presence of voltage-dependent loads.
$\bullet$ Analysing the effects of LAAs launched at different nodes of the DN and obtaining the minimum number of IoT-controllable devices required for a detrimental attack.
$\bullet$ Proposing a novel sequential game to counter the effect of LAAs using grid reconfiguration, considering the potential errors and uncertainties associated with localizing the attack.
$\bullet$ Using BO to obtain the game’s Stackelberg equilibrium to reduce its computational complexity.
The rest of the paper is structured as follows: Section II presents the implemented models and definitions in the paper. Section III provides an analytical study on the effects of LAAs on distribution systems. Subsequently, IV presents the proposed LAA mitigation scheme. Section V outlines the obtained numerical results and provides discussions. Finally, Section VI concludes the paper.
Figure 1: Summary of the proposed attacker-defender interaction.
II Preliminaries
In this section, we introduce the system load and power flow models for the DNs considered in this research.
II-A DN Model
The DN under study is represented by the connected directed graph $G=\{\mathcal{N},\mathcal{L}\}$, where $\mathcal{N}=\{1,2,\ldots,N\}$ denotes the set of buses and $\mathcal{L}$ denotes the set of branches. This graph has a radial structure, hence it is a tree. Except for bus $1$, which is the root, each bus is referred to as the ‘child’ of its parent bus which is the adjacent bus closer to bus $1$ by one branch. Thus, the set of branches is defined as =$\mathcal{L}=\{(\pi_{i},i)\ |\ \pi_{i},\ i\in\mathcal{N}\}$, where $\pi_{i}$ represents the parent of bus $i$. In this configuration, bus $1$ represents the generator bus. We denote by $\mathcal{D}_{k}$ the set of buses which forms the unique path connecting bus $1$ to bus $k$, excluding bus $1$ and including bus $k$. The depth of each bus represents the distance in terms of the number of branches between that bus and the root bus.
II-B Load Model
This subsection introduces the load models which are implemented in the rest of the paper.
II-B1 ZIP Load Model
ZIP and exponential load models are the most implemented ones in the industry [29]. Experimental values for voltage dependency of loads are captured and fitted in the ZIP model in [30], which we integrate into our formulations. The power demand under the ZIP load model is given in [28] as follows:
$s^{zip}_{i}(v_{i})=p^{l}_{i}(\alpha_{p}+\beta_{p}v_{i}+\gamma_{p}v^{2}_{i})+jq^{l}_{i}(\alpha_{q}+\beta_{q}v_{i}+\gamma_{q}v^{2}_{i}),$ (1)
where $\alpha_{k}+\beta_{k}+\gamma_{k}=1$, where $k=\{p,q\}.$ Furthermore, $v_{i}$ is the nodal voltage while $p^{l}_{i}$ and $q^{l}_{i}$ are active and reactive load demands of bus $i$ at rated voltage. The ZIP load model captures the voltage dependency of real-world loads.
II-B2 ZP Approximation
Based on (1), the ZIP model is a function of both $v_{i}$ and $v^{2}_{i}$. This causes the optimization tasks involving the power flow in the presence of ZIP loads to become nonconvex and complex. To tackle this problem, [31] have provided an approximate model for ZIP loads given by
$s^{zp}_{i}(v_{i})=p^{l}_{i}(\alpha^{\prime}_{p}+\gamma^{\prime}_{p}v_{i}^{2})+jq^{l}_{i}(\alpha^{\prime}_{q}+\gamma^{\prime}_{q}v^{2}_{i}),$ (2)
where $\alpha^{\prime}_{p}=\alpha_{p}+\frac{\beta_{p}}{2}$ ,$\alpha^{\prime}_{q}=\alpha_{q}+\frac{\beta_{q}}{2}$ ,$\gamma^{\prime}_{p}=\gamma_{p}+\frac{\beta_{p}}{2}$ , and $\alpha^{\prime}_{q}=\gamma_{q}+\frac{\beta_{q}}{2}$, while $s^{zp}_{i}(v_{i})=p_{i}^{zp}(v_{i})+jq_{i}^{zp}(v_{i})$. The new coefficients in the ZP model are obtained by the binomial approximation method. The ZP approximation is valid as long as the voltage is close enough to the nominal value i.e., while $|v_{i}-v_{nom}|\leq 0.1$, the ZP approximation is valid [31] ($v_{nom}$ is the nominal voltage).
II-C Power Flow Equations
II-C1 Branch Flow Model
The branch flow model encapsulates the complete AC power flow, with the equations describing the system state as follows [32]:
$\sum_{k:i\rightarrow k}s_{i,k}=s_{\pi_{i},i}-z_{\pi_{i},i}\ |I_{\pi_{i},i}|^{2}-s_{i},$ (3)
where $v_{\pi_{i}}-v_{i}=z_{\pi_{i},i}\ I_{\pi_{i},i},$ $s_{\pi_{i},i}=v_{\pi_{i}}\ I_{\pi_{i},i}^{*}$. Note that superscript $(\cdot)^{*}$ denotes the conjugate of a complex number and $\pi_{i}$ represents the parent node of the bus $i$. Furthermore, terms $s_{i,k}$ and $I_{\pi_{i},i}$ illustrate the apparent power and current running between two nodes, $z_{\pi_{i},i}$ is the impedance of the line, and $s_{i}$ is the apparent power in bus $i$.
II-C2 Linearized Distribution Flow
Linearized distribution flow (LinDistFlow) simplifies the branch flow model described in (3) by neglecting branch power losses [33] and is widely adopted in several DN studies. The power flow equations under this model are given by $\sum_{k:(i,k)\in\mathcal{L}}p_{ik}=p_{ji}-p_{i}$, $\sum_{k:(i,k)\in\mathcal{L}}q_{ik}=q_{ji}-q_{i}$, as active and reactive power balances and $v^{2}_{i}=v_{j}^{2}-2(r_{ij}p_{ji}+x_{ij}q_{ji})$ the equation for finding subsequent voltage profile. Note that $p_{ik}$ and $q_{ik}$ are active and reactive power running the corresponding line, $p_{i}$ and $q_{i}$ are active and reactive loads in bus $i$, while $r_{ij}$ and $x_{ij}$ are resistance and reactance of lines.
II-D Reconfiguration of DN
Network reconfiguration involves modifying the DN’s topology by adjusting the open/closed state of its switches [22]. In this subsection, a set of MILP optimization constraints is introduced primarily to determine the configuration that maintains the nodal voltages closest to their nominal values. We modify the formulation in [34] to accommodate the ZP approximation of the loads to capture its voltage dependency. We also change the power flow model to the LinDistFlow. As we show in Section V, these approximations [31, 33] result in a MILP the network reconfiguration problem and provide significant computational speed-ups compared to the MISOCP.
$\bullet$ Connectivity Constraints: This set of constraints determines the connection between the nodes while keeping the overall graph radial. For brevity, we have omitted these constraints here, and they can be found in [34].
$\bullet$ Power Flow Constraints: Below, we present the optimization problem’s power flow constraints, which are taken from the DistFlow and ZP models:
$|p_{ij}|,|q_{ij}|\leq Mb_{ij},$ (4)
$\sum_{j:(i,j)\in\mathcal{L}}p_{ij}=p_{i}^{f},\sum_{j:(i,j)\in\mathcal{L}}q_{ij}=q_{i}^{f},\hskip 28.45274pti\in\mathcal{N}^{f}$ (5)
$\sum_{j\in\mathcal{N}}p_{ji}-p_{ij}=p_{i}^{ZP},\sum_{j\in\mathcal{N}}q_{ji}-q_{ij}=q_{i}^{ZP},\hskip 8.53581pti\in\mathcal{N}\setminus\mathcal{N}^{f}$ (6)
$s^{zp}_{i}(v_{i})=p^{l}_{i}(\alpha^{\prime}_{p}+\gamma^{\prime}_{p}v_{i}^{2})+jq^{l}_{i}(\alpha^{\prime}_{q}+\gamma^{\prime}_{q}v^{2}_{i}),$ (7)
$\underline{v_{i}}^{2}\leq v_{i}^{2}\leq\overline{v_{i}}^{2},$ (8)
$\hat{v}^{2}_{ij}\leq Mb_{ij},$ (9)
$\hat{v}^{2}_{ij}\leq v^{2}_{i}-2(r_{ij}p_{ij}+x_{ij}q_{ij}),$ (10)
$v^{2}_{i}=\sum_{j\in\mathcal{N}}\hat{v}^{2}_{ji}.\hskip 28.45274pti\in\mathcal{N}\setminus\mathcal{N}^{f}$ (11)
Here $b_{ij}$ represents the elements of the connectivity matrix of the grid, $\mathcal{N}^{f}$ is the set of substation bus(es), $M$ is the disjunctive parameter, $p_{i}^{f}$ and $q_{i}^{f}$ are active and reactive power flows from substation bus $i$, $\underline{v_{i}}$ and $\overline{v_{i}}$ are voltage limits, and $\hat{v}$ is an auxiliary variable. Equations (4) - (6) represent the power flow constraints, (7) is ZIP load constraint, and (8) - (11) are voltage constraints. The auxiliary variable of $\hat{v}$ makes the optimization follow disciplined convex programming rules in Python. Additionally, we consider all the lines to have switches.
II-E Load-Altering Attack Model
Some major manufacturers of high-wattage IoT-controllable devices have acknowledged the presence of security vulnerabilities in their products [11]. The concept presented in the LAAs involves attackers leveraging these vulnerabilities to transform a group of such devices into bots and toggle them on and off. This coordinated action is intended to disrupt the stability of the system. Based on this, to power balance equations change into
$\begin{cases}\sum_{j\in\mathcal{N}}x_{ji}-x_{ij}=x_{i}\hskip 36.98866pti\in\mathcal{N}\setminus\mathcal{N}^{f},i\notin\mathcal{N}^{a},\\ \sum_{j\in\mathcal{N}}x_{ji}-x_{ij}=x_{i}+x^{a}\hskip 11.38092pti\in\mathcal{N}\setminus\mathcal{N}^{f},i\in\mathcal{N}^{a},\end{cases}$ (12)
where $x$ is a short-hand notation, with $x=p$ for active power and $x=q$ for reactive power. Herein, $p^{a}$ and $q^{a}$ are additional active and reactive powers change due to the LAA in bus $i$ and $\mathcal{N}^{a}$ is the set of attacked buses. In the following, based on the provided models and context, we first analyze the impact of LAA on DNs. Then, we propose a game-theoretic mitigation scheme for these attacks.
III Effects of LAA on DNs
In this section, we analyze the impact of LAA on DNs. Our objective is to derive closed-form expressions for the voltage profile of the network with voltage-dependant loads under LAA and for the minimum compromised devices required to cause nodal voltage safety violations. Note that grid’s voltage under LAA can also be computed by solving the power flow equation (3) through an iterative approach such as the backward-forward sweep (BFS) technique. However, unlike closed-form equations (which we derive in this section), the iterative techniques does not yield analytical insights into the impact of LAAs. Furthermore, the closed-form expressions obtained in this section are essential in the design of the defense strategies to mitigate LAAs.
III-A Closed-form Approximation of Nodal Voltages
To derive the closed-form expressions for the system voltages under LAAs, we make two approximations: (i) employing LinDistFlow formulations and (ii) utilizing the ZP model.
III-A1 Without LAA
First, we model the DN without LAA and analyze the power flow equations. Integrating (2) into voltage equation results in
$v_{k}=\sqrt{v_{1}^{2}-2\sum_{i\in\mathcal{D}_{k}}\left(r_{\pi_{i},i}p^{zp}_{\pi_{i},i}+x_{\pi_{i},i}q^{zp}_{\pi_{i},i}\right)},$ (13)
where $p_{\pi_{i},i}^{zp}=p_{\pi_{i},i}(\alpha^{\prime}_{p}+\gamma^{\prime}_{p}v_{i}^{2}),$ $q_{\pi_{i},i}^{zp}=q_{\pi_{i},i}(\alpha^{\prime}_{q}+\gamma^{\prime}_{q}v_{i}^{2}).$ Next, we perform a variable change $(u_{k}=v_{k}^{2}),$ which results in a set of linear equations, which can be written in matrix form as $\textbf{U}_{(N-1)\times 1}=\Omega_{(N-1)\times N}\begin{bmatrix}1\\ \textbf{U}\end{bmatrix}_{N\times 1}$, where U is the vector of squares of voltages, and $\Omega_{(n-1)\times n}$ is the matrix with entries:
$\omega_{i,1}=1-\sum_{m\in\mathcal{D}_{i}}(2r_{\pi_{m},m}p_{\pi_{m},m}^{0}\alpha^{\prime}_{p}+2x_{\pi_{m},m}q_{\pi_{m},m}^{0}\alpha^{\prime}_{q}),$ (14)
[…]
where $2\leq i,k\leq N$ ($N$ is number of buses). We re-write the system of linear equations as $(\textbf{I}_{(N-1)\times(N-1)}-\Omega^{\prime}_{(N-1)\times(N-1)})\ \textbf{U}_{(N-1)\times 1}=\Omega^{\prime\prime}_{(N-1)\times 1}$, in which $\Omega^{\prime\prime}_{(N-1)\times 1}=[\omega_{2,k}]$, $\Omega^{\prime}_{(N-1)\times(N-1)}=[\omega_{i,k}]$ for $i=\{3,4,...,N\}$, and $k\in\mathcal{N}$.
III-A2 With LAA
Here, we analyze the voltage profile of the network under LAA. For this, we integrate the introduced LAA in Section II-E into (13) and obtain
$v_{k}^{a}=\sqrt{v_{1}^{2}-2\Delta_{k}-2p^{A}_{a}r_{k,a}-2q^{A}_{a}x_{k,a}},$ (16)
in which $\Delta_{k}=\sum_{i\in\mathcal{D}_{k}}\left(r_{\pi_{i},i}p^{zp}_{\pi_{i}^{,}i}+x_{\pi_{i},i}q^{zp}_{\pi_{i},i}\right)$,
$r_{k,a}=\sum_{i\in\{\mathcal{D}_{a}\cap\mathcal{D}_{k}\}}r_{\pi_{i},i},$ and $x_{k,a}=\sum_{i\in\{\mathcal{D}_{a}\cap\mathcal{D}_{k}\}}x_{\pi_{i},i}$. This change results in a new set of coefficient matrices. To calculate the attacked system’s square of voltages vector $(\textbf{U}^{A})$, we solve $\textbf{U}^{A}_{(N-1)\times 1}=\Omega^{A}_{(N-1)\times N}\begin{bmatrix}1\\ \textbf{U}^{A}\end{bmatrix}_{N\times 1}$. To obtain the coefficients matrices, (12) is dragged into LinDistFlow as the ZP model is imposed on them. The final results are $\omega_{i,k}^{A}=\omega_{i,k}+\omega_{i,k}^{a}$, for $i\geq 2$ and $k\geq 1$; $\Omega^{A}_{(N-1)\times N}=[\Omega^{A^{\prime\prime}}_{(N-1)\times 1}\ \Omega^{A^{\prime}}_{(N-1)\times(N-1)}]$, in which for $i\geq 2$ and $k\geq 2$ and $\omega_{i,1}^{a}=\sum_{c\in\{\mathcal{D}_{i}\cap\mathcal{D}_{a}\}}-2p_{a}^{A_{0}}\alpha^{\prime}_{p}r_{\pi_{c},c}-2q_{a}^{A_{0}}\alpha^{\prime}_{q}x_{\pi_{c},c}$ , $\omega_{i,k}^{a}=\begin{cases}-2r_{\pi_{i},i}p^{a}\gamma^{\prime}_{p}-2x_{\pi_{i},i}q^{a}\gamma^{\prime}_{q},\ \ \text{if}\hskip 5.69046pti\in\mathcal{D}_{a}\\ \omega_{\pi_{i},k}^{A},\ \ \text{otherwise}.\end{cases}$
If there is more than one attacked node at a time, the impacts will be summed up. As a result, the new resulting coefficients are: $\omega^{a_{t}}_{i,j}=\sum_{a\in\mathcal{A}}\omega^{a}_{i,j},\ \forall i\in[2,N]\ \&\ j\in\mathcal{N}$. Subsequently, $\omega_{i,k}^{A}=\omega_{i,k}+\omega_{i,k}^{a_{t}}$, for $i\geq 2$ and $k\geq 1$.
Intuitively, higher power flows in distribution lines cause bigger voltage drops. The additional demand caused by LAAs in deeper nodes leads to higher load flows in more lines (to convey the power from the root to the LAA location). As a result, these LAAs cause more drastic voltage drops.
III-B Analytical Insights into the Attack Impact
Next, we use the closed-form expressions obtained above to derive analytical insights into the attack impact. First, note that when there is an attack in a leaf bus (the last bus of each branch), $r_{k,a}$ and $x_{k,a}$ have the highest possible values. As a result, the voltage drop resulted from $(-2p^{a}r_{k,a}-2q^{a}x_{k,a})$ in (16) is higher and obtained voltages shrink. In conclusion, attacks on the leaf buses yield the most detrimental effects.
While this result is somewhat expected for DNs, we further obtain the minimum number of attacked devices, which leads to voltage safety violations. We call such a threat the “critical attack”. For this, we consider the voltage of the leaf bus as a known variable ($v_{th}$). The new unknown variable is $p^{a}$, and based on the attacked device type, we can find $q^{a}$ via $q^{a}=\frac{q_{d}}{p_{d}}\ p^{a}$. So the new set of coefficients for obtaining voltages of buses except for the leaf one and the active power of the critical attack is forming $\Omega^{d}$ ($\Omega^{d}=[\Omega^{d^{\prime\prime}}\Omega^{d^{\prime}}]$) in which
$\omega_{i,1}^{d}$ =$=\begin{cases}\omega_{i,1}+v_{th}^{2}\omega_{i,a},\ \ \text{if}\ i\neq a,\\ \omega_{i,1}+v_{th}^{2}(\omega_{i,a}-1),\ \ \text{if}\ i=a,\end{cases}$ (17)
$\omega_{i,k}^{d}$ =$=\begin{cases}\sum_{c\in\mathcal{D}_{i}}-2r_{\pi_{c},c}\alpha^{\prime}_{p}-2\frac{q_{d}}{p_{d}}x_{\pi_{c},c}\alpha^{\prime}_{q},\ \ \\ \quad\quad\quad\ \text{if}\ k=a,i\in\mathcal{D}_{a},\\ \omega_{\pi_{i},k}^{d},\ \ \text{if}\ k=a,i\notin\mathcal{D}_{a},\\ \omega_{i,k},\ \ \text{otherwise}.\end{cases}$ (18)
Hence, we can solve the linear system of equations given by $(\textbf{I}^{a}-\Omega^{d^{\prime}})\textbf{X}=\Omega^{d^{\prime\prime}}$, in which X is a vector with the same dimension as U where all elements of it are the same as U except for one in which X contains $p^{a}$ instead of $u_{a}$ (since we already know $u_{a}=v_{th}^{2}$). Additionally, $\textbf{I}^{a}$ is the identity matrix except for the element $(a,a)$ which equals 00. This gives the bus voltages as well as the required $p^{a}$. Then, we can find the number of required devices in bus $n$ using $c(n)=\frac{p^{a}}{p_{d}}$.
The closed-form expressions provided in this section are not only used for identifying the worst effect of the LAAs, but also in the following sections, they inspire our mitigation method to find the optimal defensive action. Additionally, the introduced “critical attack” has been implemented in finding the optimal action of the strategic attacker later in this paper.
IV Mitigating LAA via Reconfiguration
In this section, we introduce a novel technique to mitigate LAAs by reconfiguring the DN topology. We exploit a sequential game-theoretic interaction in which, following the LAA launched by the attacker, the defender reconfigures the network to react optimally to the threat.
IV-A Mitigation Design and Intuition
The intuition behind the proposed defense technique of reconfiguring the DN lies in the analytical insights derived in Section III. Based on our analysis, LAAs targeting the leaf buses of the DN lead to the greatest attack impact (i.e. deviation of the voltage from the nominal values). In this context, reconfiguring the DN, changes the position of the leaf buses, thus alleviating the attack impact. It is worth noting that the proposed mitigation leverages the pre-existing capabilities of the DN (e.g., devices enabling network reconfiguration are primarily installed to reduce power losses and/or voltage deviations) and, hence, does not require new infrastructure.
Furthermore, in the proposed scheme, the system will be reconfigured only when an attack is detected, thus avoiding unnecessary reconfigurations to mitigate attacks (note that cyberattacks are somewhat rare events). The attack detection module can be based on existing model-driven or data-driven approaches for detecting LAAs. The reader can refer to past works, including [26] and [27] in this area for more details.
IV-B Stackelberg Game for Attack Mitigation
We model the strategic interaction of the attacker and the defender via a non-cooperative Stackelberg game in which the attacker first chooses a bus to launch an LAA, and subsequently, the defender exploits the flexibility of the system to reconfigure it and mitigate the attack. This sequential game models the reactive mitigation framework. Table II summarizes the assumptions and details of the game.
TABLE II: Non-cooperative game assumptions.
Player Attacker Defender
Role Leader Follower
Knowledge Potential defensive reactions Grid’s topology
Actions Switching loads Changing the topology
Goal Violate voltage const. Min. attack impact
A Stackelberg game with two players consists of a leader and a follower. The leader always commits their action first to maximize their own objective function by anticipating the follower’s reaction. Then, given the action of the leader, the follower picks their best (i.e., optimal) response to maximize their own objective function. Since LAAs are rare incidents in the network, we propose a reactive mitigation method instead of a preventive one. The proposed Stackelberg game can be defined as =$\mathcal{H}=\{(A,D),(\mathcal{S}_{A},\mathcal{S}_{D}),(F_{A},F_{D})\}$, in which $A$ and $D$ are the players (attacker and defender), $\mathcal{S}_{A},\mathcal{S}_{D}$ denote sets of actions of players, and $F_{A},F_{D}$ denote their rewards.
The attacker’s set of actions, $\mathcal{S}_{A}$, is launching an LAA in any of the buses (one or two buses at a time by assumption). The set of defense actions is the set of all possible system reconfigurations discussed in Section II-D such that $\mathcal{S}_{D}=\{1,2,\ldots,N_{B}\}$ denotes the set of indices of all possible reconfiguration matrices $B=[b_{ij}]_{1\leq i,j\leq N}\in\{0,1\}^{N\times N}$ whose entries meet the connectivity constraints and $N_{B}$ represents the number of such matrices. Henceforth, we denote by $B(d)=[b_{ij}(d)]_{1\leq i,j\leq N}$ the adjacency matrix of the network as a result of a specific defense $d\in\mathcal{S}_{D}$. The attacker aims to maximize the voltage deviation, as a result, we define $F_{A}(d,a)=\sum_{n\in\mathcal{N}}\left|v^{2}_{nom}-v^{2}_{n}(d,a)\right|$, as the attacker’s objective function.
The defender’s objective is to reconfigure the system to minimize the square of the voltage deviation above. Achieving this goal with minimal changes can decrease the maintenance requirements for switches and reduce the likelihood of switching failures. To take this factor into account, we add a penalty term and the resulting reward function of the defender is
$F_{D}^{perf}=-\sum_{n\in\mathcal{N}}\left|v^{2}_{nom}-v^{2}_{n}(d,a)\right|-pen(d),$ (19)
where $pen(d)=\sum_{i=1}^{N}\sum_{j=1}^{N}\left|b^{pre}_{ij}-b_{ij}(d)\right|$ is the penalty term for enforcing a system reconfiguration with the minimum switches possible. Note that $b^{pre}_{ij}$ is the component of the grid’s connectivity matrix before reconfiguration. $F_{D}^{perf}$ is relevant when the defender is capable of perfect attack localization. However, due to noisy measurements, the defender might not be able to do this. Instead, we assume that the defender is only able to locate a neighbourhood of the attack, i.e., a connected cluster of buses which contains the bus under attack. We further assume that the defender has a favourite candidate bus under attack, denoted by $n_{att}$, but does not discard the attack possibilities of other buses in the located cluster. To model this, we define a discrete probability vector $\sigma=[\sigma_{1},\sigma_{2},\ldots,\sigma_{N}]$ with entries:
$\sigma_{\ell}=\begin{cases}\rho,&\textbf{if}\hskip 8.5359pt\ell=n_{att},\\ \frac{1-\rho}{|\mathcal{A}_{n_{att}}|},&\textbf{if}\hskip 8.5359pt\ell\in\ \mathcal{A}_{n_{att}},\\ 0,&\textbf{otherwise}.\end{cases}$ (20)
Above, $\sigma_{\ell}$ represents the likelihood that the defender assigns to bus $\ell$ being under attack such that $\sigma_{\ell}\in[0,1]$ and $\sum_{\ell\in\mathcal{N}}\sigma_{\ell}=1$. Additionally, $\rho\in[0.5,1]$ denotes the likelihood of the defender’s favourite candidate. The subset $\mathcal{A}_{n_{att}}$ is the set of buses directly adjacent to $n_{att}$, along with the buses directly adjacent to those; all these buses are considered as the other potential candidates by the defender. Their likelihood is the remaining probability $1-\rho$ split equally between the $|\mathcal{A}_{n_{att}}|$ other candidate buses.
Taking this uncertainty (of detecting precisely the attack location) into account at the defender’s end results in the following reward:
$F_{D}(d,a)=-\sum_{\ell\in\mathcal{N}}\sigma_{\ell}\sum_{n\in\mathcal{N}}\left|v^{2}_{nom}-v^{2}_{n}(d,a_{\ell})\right|-pen(d),$ (21)
which represents the expected reward over this uncertainty. Obtaining the optimal attack and defense requires computing the Stackelberg equilibrium. As discussed earlier, in this paper, the attacker commits the attack first and then the defender reacts.
Definition 1.
The best response of the defender to an action $a\in\mathcal{S}_{A}$ is defined as:
$r(a)=\arg\max_{d\in\mathcal{S}_{D}}F_{D}(d,a).$ (22)
Definition 2.
A profile of actions $(a^{*},d^{*})\in(\mathcal{S}_{A},\mathcal{S}_{D})$ is a Stackelberg equilibrium iff
$\begin{cases}F_{A}(r(a^{*}),a^{*})&\geq F_{A}(r(a),a),\hskip 19.91684pt\forall a\in\mathcal{S}_{A}\\ \hskip 51.21504ptd^{*}&=r(a^{*}).\end{cases}$ (23)
Intuitively, the attacking action at the Stackelberg equilibrium is the one maximizing the attacker’s reward under the defender’s best reaction. Furthermore, the defender’s best reaction to $a^{*}$ is its Stackelberg equilibrium action. A Stackelberg equilibrium is ensured to exist if the defender’s optimal response exists for every attack. Assuming that normally open points exist in the distribution system, this ensures that at least one system reconfiguration is possible and that the discrete feasible set in (22) is non-void, leading to the existence of the solution (if the corresponding constraints are met).
IV-C Bayesian Optimization
BO provides algorithms for optimizing black-box functions, whose mathematical expression is unknown or too complex to analyze [35]. Instead, BO relies on the function’s observed values for given inputs. This makes BO also suitable for optimizing computationally expensive functions to evaluate exhaustively [36]. In our case, obtaining the best response function for all possible attacks requires solving a separate optimization problem for each attack, leading to significant computational complexity. To address this issue, we consider the attacker’s reward to be a black-box function and utilize BO to explore and approximate the optimal attack efficiently. Similarly to [37], we first build a probabilistic model ($F^{p}_{A}$) for $F_{A}(a,r(a))$ using a Gaussian process $(\mathcal{GP})$. This model is based on sample attacks $\mathcal{T}=\{a^{s}_{i}\}$ and their corresponding $\{F_{A}(a^{s}_{i},r(a^{s}_{i}))\}$, which is obtained using:
$F^{p}_{A}\sim\mathcal{GP}(m(a),k(a,a^{\prime})),$ (24)
where $(m(a))$ is the mean function and $(k(a,a^{\prime}))$ is the covariance kernel. Note that we sample the attack locations to form the $\mathcal{T}$ from all branches to have a good initial estimate model. Afterwards, we use the expected improvement (EI) function to pick the next attack action for evaluation via:
$a^{\mathrm{next}}=\arg\max_{a\in\mathcal{S}_{A}}EI(a),$ (25)
and add this new point to the $\mathcal{T}$ and update $F^{p}_{A}$. The EI function is $EI=\mathbb{E}[\max\{0,F_{A}^{\max}-F^{p}_{A}(a)\}]$. The process is repeated until the stopping criteria are met.
Algorithm 1 describes the method of finding the players’ actions at the Stackelberg equilibrium.
1
Data: $\mathcal{H}\ ,\ \sigma=\{\sigma_{1},...,\sigma_{N}\}$
Result: $r(a^{*})$ , $a^{*}$
1:  Sample $\{a^{i}\}_{i=1}^{n_{0}}$ from $\mathcal{S}_{A}$, form the set $\mathcal{T}$, and compute $r(a),\ \forall a\in\mathcal{T}$ by solving       $r(a)=\arg\max_{d\in\mathcal{S}_{D}}F_{D}(d,a)$ s.t. connectivity and power flow constraints as an MILP optimization problem via SCIPY and CVXPY;
22:  Compute $F_{A}(r(a),a),\ \forall a\in\mathcal{T}$;
33:  Construct a probabilistic model of the $F_{A}(r(a),a)$ as in (24);
44:  Find the next query point $(a^{\mathrm{next}})$ using (25), attach it to $\mathcal{T}$, and compute $F_{A}(r(a^{\mathrm{next}}),a^{\mathrm{next}})$;
55:  Repeat until the stop condition is met;
66:  Choose $a^{*}=\arg\max_{a\in\mathcal{T}}F_{A}(r(a),a)$;
77:  Compute $r(a^{*})=\arg\max_{d\in\mathcal{S}_{D}}F_{D}(d,a^{*})$ as in step 1;
Algorithm 1 Computing the Stackelberg equilibrium using BO.
The constraints of the reconfiguration optimization (as in steps one and seven of the algorithm 1) are linear in the square of the voltages, hence, justifying our choice of the distance between the squares of voltages in the reward function, leading to a linear program instead of a quadratic one (obtained by a variable change $u_{i}=v_{i}^{2}$). The resulting MILP optimizations are carried out in Python, the optimization modeling language is CVXPY, and the solver is SCIPY. We find the reward of the attacker for $(r(a),a),\ \forall a\in\mathcal{T}$. Finally, the action corresponding to the maximum of $F_{A}(r(a),a)$ is selected as the attacking strategy in our Stackelberg formulation. Additionally, $r(a^{*})$ corresponding to the Stackelberg attack action $a^{*}$ is the optimum defensive strategy at the Stackelberg equilibrium: $d^{*}=r(a^{*})$.
IV-D Resource-Constrained Attacker
Drawing from our discussions in Section III-B, each bus has a distinct “critical attack” leading to voltage constraint violation. Given the attacker’s tendency for launching such a “critical attack”, their potential action will not only occur across different buses but also vary in magnitude.
To accommodate this feature, we define a new game =$\mathcal{H}^{\prime}=\{(A,D),(\mathcal{S}_{A},\mathcal{S}_{D}),(F^{\prime}_{A},F_{D})\}$ in which the attacker launches the “critical attack” which we call them the “resource-constrained attacker”. Note that the critical attack in each bus is manipulating the minimum devices ($c(n)$) in that bus to cause a voltage constraint violation. Indeed, $c(n)$ for each attack is unique and varies with the attack location. In this modified game, $F^{\prime}_{A}$ comprises two components: the total nodal voltage deviation and the attack magnitude. The attacker seeks to maximize the former while minimizing the latter. However, these two terms cannot be simply summed due to their disparate physical characteristics. Therefore, we propose the following reward:
$F^{\prime}_{A}(d,a)=(1-\lambda)F_{A}^{\mathrm{norm}}(d,a)-\lambda c^{\mathrm{norm}}(a),$ (26)
in which $0\leq\lambda\leq 1$, $F_{A}^{\mathrm{norm}}(d,a)=\frac{F_{A}(d,a)}{\sum_{i\in\mathcal{S}_{A}}F_{A}(d,i)}$, and $c^{\mathrm{norm}}(a)=\frac{c(a)}{\sum_{i\in\mathcal{N}^{L}}c(i)}$. The parameter $\lambda$ trades off between the two components of the objective function. If $\lambda=0$, the attacker only cares about maximizing the harm caused in the voltage profile; and if $\lambda=1$ the attacker only cares about minimizing the attacked devices. The rest of the components of $\mathcal{H}^{\prime}$ are the same as $\mathcal{H}$. The process of computing the Stackelberg equilibrium is similar to $\mathcal{H}$, and we only need to plug in the attacker’s new objective function $F^{\prime}(d,a)$.
V Results and Discussions
Our simulations are conducted using the IEEE 33-bus and 69-bus systems. There are 32 and 48 load buses in the test grids, respectively, which could be attacked. The ZIP load coefficients are set to the residential load-type F introduced in [30].
V-A Critical Attack
Here, we conduct spatial analyses to determine the most effective location for launching LAAs. To quantify this, we use the load profile obtained from [38] for the date 05/05/2024 as the base load (without LAAs), which contains the hourly load demand in New York, US. To mimic this load profile in the 33-bus grid, we project the ratio of load changes at different hours onto the nominal load of the test network. Table III presents the number of devices required for the critical attack in three of the leaf buses of the test case during the different hours of the day. These numbers are computed via the equations in Section III-B. We can see that the attack on the deepest bus requires fewer devices to be manipulated. This conclusion confirms our insights from Section III-A. Furthermore, it also shows the dependency on the type of load (air conditioner, resistive load, etc.) and the associated ZIP load coefficients.
TABLE III: Numbers of compromised devices required to cause voltage safety violations in the 33-bus grid during different hours.
Time Att. location (bus) 03:00 (Least load) 09:00 18:00 (Peak load)
Air Conditioner 18 603 459 38
25 5538 5168 612
33 1498 1079 248
Resistive heater 18 256 134 17
25 2573 1582 186
32 612 441 85
Since the results in Table III are obtained by the approximation discussed in Section III, we evaluate how effective these attacks are when considering the full AC power flow model. To evaluate the extent of the errors (between the voltages computed using the analytical results and the full AC power flow model), we compare the obtained voltage profile with the results of BFS. Fig. 2 shows the voltage profile calculated by the two methods during the peak load demand and the corresponding critical attack. We remark that the actual errors of computing the nodal voltages with our approximations never exceed $1\%$, thus proving the validity of the analytical results.
Figure 2: Voltage profile of the attacked (on Bus 18) 33-bus test case with the proposed closed-form equations and the accurate model.
V-B LAA Mitigation
Next, we examine the proposed LAA mitigation method. The base load profiles (without LAAs) of 33-bus and 69-bus test cases are $60\%$ and $30\%$ of their nominal load profiles in MATPOWER. For scenarios i) and ii), we consider LAA attacks of magnitude $p_{i}^{a}=q_{i}^{a}=0.30\ p.u.$ in which $i$ is the index of the attacked bus. This magnitude of the attack is significant enough to cause voltage safety violations in all the load buses (except those adjacent to the root) and, hence, needs to be mitigated. In 33-bus grid, {(8-21), (9-15), (12-22), (18-33), (25-29)} and in 69-bus grid {(11-43), (13-21), (15-46), (27-65), (50-59)} are the normally open lines which can be closed during the system reconfiguration. Four scenarios are considered next: i) accurate attack localization, ii) errors in attack localization, iii) errors in attack localization and resource-constrained attacker, and iv) attacks on two nodes.
i) Accurate Attack Localization: Here, we set $\sigma_{n_{att}}=1$ and $\sigma_{\ell}=0\hskip 4.26773pt\forall\hskip 4.26773pt\ell\in\mathcal{N}^{L}\setminus\mathcal{N}^{a}$. Table IV presents the players’ actions at the Stackelberg equilibrium. Based on these results, in the 33-bus grid, the attack is launched on Bus 33. Fig. 3 shows the voltage profiles of the grid under attack before and after reconfiguration. The results show that the proposed mitigation method is able to return the voltage profile within the constraints, hence mitigating the effects of LAA. Although the LAA on Bus 18 causes the greatest impact in terms of voltage deviations, a strategic attacker who can anticipate the defender’s reaction chooses to launch the attack on Bus 33 instead to maximize their payoff. Furthermore, in the 69-bus grid, the attack is launched on Bus 27 and lines (50-59) and (58-59) change their state to reconfigure the network.
Figure 3: Voltage profile of the attacked 33-bus grid before and after mitigation (reconfiguration).
ii) Errors in Attack Localization: In this scenario, following the worst case accuracy of the detection algorithm in [27], we consider $\rho=0.7$ ($70\%$ of certainty about the location of the attack), and then the remaining $30\%$ is split equally between other buses in the neighbourhood (see Section IV). The results for this scenario are presented in Table IV. In this scenario, the defense action should keep the voltage within the desired constraints, assuming an LAA in any of the candidate buses. This uncertainty causes a system reconfiguration that necessitates more switching. Fig. 4 represents the voltage profile of the 33-bus grid after the reconfiguration if any of the three suspect buses are attacked. We can notice that the attack will be successfully mitigated and the voltage constraint will not be violated in any case.
Note that the players’ actions for the 69-bus system are the same as in scenario i).
Figure 4: Voltage profiles of 33-bus grid attacked at any of the suspect buses after mitigation.
iii) Resource-Constrained Attacker: Here, we present the resource-constrained attacker introduced in Section IV-D with $\lambda=0.5$ in their objective function. Similar to the two previous cases, the results of this scenario can also be found in Table IV. We observe that in this scenario, the attacker chooses to attack Bus 18, as the impact of load alteration on this bus will be the greatest (as the number of compromised devices is taken explicitly into account). The 69-bus case maintains the same Stackelberg equilibrium actions as the past two scenarios.
iv) Attacks on two nodes: In this scenario, with insufficient vulnerable loads, the attacker targets two locations. In the 33-bus grid, this results in affected buses 32 and 33, with a defensive response similar to Scenario ii). In the 69-bus grid, buses 26 and 27 are targeted, showing the same defensive pattern as before. Table V represents the total voltage deviations ($\sum_{i\in\mathcal{N}}\left|v_{nom}-v_{i}\right|$) in the grid for different scenarios. Note that, in scenario ii), the voltage deviation of the 33-bus grid drops, but the defender needs to commit more switching, which is not desirable. Furthermore, a portion of the reduced voltage deviation in scenario iii) should be attributed to the smaller attack launched by the resource-constrained attacker.
TABLE IV: Attacked bus(es) and closed/open lines at the Stackelberg equilibrium of each scenario.
Scenario Attacked bus Closed line(s) Opened line(s)
33-bus 69-bus 33-bus 69-bus 33-bus 69-bus
i 33 27 (25-29) (50-59) (28-29) (58-59)
ii 33 27 (22-12), (25-29) (50-59) (11-12), (28-29) (58-59)
iii 18 27 (22-12) (50-59) (11-12) (58-59)
iv 32&33 26&27 (22-12), (25-29) (50-59) (11-12), (28-29) (58-59)
TABLE V: Total voltage deviations and switching required under each scenario.
Scenario Number of switching Total voltage deviation (p.u.)
33-bus 69-bus 33-bus 69-bus
i 2 2 0.84 1.31
ii 4 2 0.66 1.31
iii 2 2 0.76 1.31
iv 4 2 0.65 1.30
V-C Significance of Game-Theoretic Approach
We also consider a non-strategic attacker that does not anticipate the defender’s actions. In this case, first, the attacker launches the attack, which maximizes the total voltage deviation, regardless of potential defensive reactions. Subsequently, the defender solves the optimal reconfiguration problem to mitigate the attack. The results of this approach for the 33-bus test case are presented in Table VI. We remark that both scenarios i) and ii) result in the same output. Compared to Table V for the strategic attacker, the defender always benefits. Indeed, in scenario i), the total voltage deviation is dropped from $0.84$ p.u. to $0.80$ p.u.; while in ii), the number of switches is reduced from $4$ to $2$. Note that mitigation with less switching is preferred (even with slightly higher voltage deviation). Note that the strategic attacker in the 69-bus grid picks the deepest bus (Bus 27) as the victim, which is similar to the non-strategic attacker’s choice.
TABLE VI: Non-strategic attacker’s preferred actions, obligated switching numbers, and total voltage deviations.
Scenario Attacked bus Total voltage deviation (p.u.) Number of switching
i 18 0.80 2
ii 18 0.80 2
To show the efficiency of the approximations introduced in Section II-D, we compare the computation time for the implemented MILP model with that of the similar MISOCP model. Exhibited in Table VII, for the 33-bus grid, a single MISOCP optimization takes $206.20$ s while one MILP optimization requires $19.47$ s, which means a reduction of over $90\%$ in calculation time, while for the bigger system, this reduction is even more significant ($96\%$). To further reduce the computational time, BO is implemented to minimize the number of optimizations required for the Stackelberg equilibrium determination. Table VIII contrasts the optimization counts with and without BO, demonstrating a significant reduction.
TABLE VII: Optimization times in seconds.
Op. type 33-bus grid 69-bus grid
MISOCP 206.20 s 3168.87 s
MILP 19.47 s 103.14 s
TABLE VIII: Number of optimizations required to compute the Stackelberg equillibrium
Scenario With BO Without BO
33-bus 69-bus 33-bus 69-bus
i, ii, iii 10 13 32 48
iv 41 69 496 1128
VI Conclusions
In this paper, we investigate the impact of LAAs on DNs. We derive a set of closed-form expressions for the power flow of DNs to determine bus voltages in the presence of voltage-dependent loads, with or without LAA. Then, we introduce a sequential game-theoretic approach to mitigate LAAs in DNs by network reconfiguration with minimum possible switching. Furthermore, we take into account the uncertainties in the attack localization by introducing a probability distribution over the potentially attacked nodes. To enhance the sustainability and computational speed of the Stackelberg equilibrium, a Bayesian optimization algorithm was implemented to reduce the computational burden. The proposed mitigation scheme is capable of keeping the voltage within the desired constraints. Building on our mitigation method, future work will investigate the cyber recovery step, which includes isolating the attack. Other interesting extensions are: dynamic LAAs, integrating distributed energy resources in the DN model, including the transient dynamics of the grid’s voltage, etc.